# Convergent Privacy Framework with Contractive GNN Layers for Multi-hop Aggregations

### v1

@@ 6.2. Privacy Auditing
In Table 4, we report the AUC (Area Under the Curve) score about the attack effectiveness, when $\mathsf{CARIBOU}$ is being attacked. AUC is a major metric to evaluate the membership inference attack (Wu et al., 2022a; He et al., 2021a). Specifically, AUC measures true positive rate against the false positive rate on various classification thresholds, and a score of $0.5$ suggesting random guessing. We found $\mathsf{CARIBOU}$ is very effective against Linkteller, by dropping the attack AUC from between 0.86 to 0.998 across all standard datasets (ϵ=infitalic-$\epsilon=\inf$) to less than 0.500 ($\epsilon$ ranges from 1 to 32). For G-MIA, its AUC on the 5 datasets are already lower than Linkteller by a notable margin when ϵ=infitalic-$\epsilon=\inf$ (between 0.567 to 0.702 for the 5 datasets), so the impact of $\mathsf{CARIBOU}$ is relative small. But we observe on Cora, $\mathsf{CARIBOU}$ is able to drop AUC from 0.645 to 0.500.
Finding 2.
@@
6.3. Ablation Study
Impact of $K$.
Both $\mathsf{CARIBOU}$ and GAP perform $K$-hop aggregations under $K$ aggregation layers. Here we evaluate the impact of $K$ on accuracy on the chain-structured datasets (Chain-S, Chain-M and Chain-L), as their classification results highly depend on long-range interactions. In Figure 8 of Appendix, we compare $\mathsf{CARIBOU}$ and GAP on varying $\epsilon$. The result of GAP is drawn with lines and the result of $\mathsf{CARIBOU}$ is illustrated with the colored boxes, because $\mathsf{CARIBOU}$ also depends on other hyper-parameters $C_{\mathsf{L}},\alpha_{1},\beta$ and we use the colored boxes to represent the interquartile range over 5 runs of their different value combinations.
In Figure 5(a), we show the result of one setting ($\epsilon=1$ on Chain-S), and $\mathsf{CARIBOU}$ achieves higher accuracy at every $K$. This observation is consistent in other settings as shown in Figure 8. On the other hand, the classification accuracy is fluctuated when $K$ increase, for both $\mathsf{CARIBOU}$ and GAP. For instance, the best accuracy of $\mathsf{CARIBOU}$ is achieved when $K$ equals to $6$ or $7$, which is close to the chain length (i.e., $8$ for Chain-S). This result suggests the added DP noises could outweigh the benefits of aggregations at some point, though $\mathsf{CARIBOU}$ handles this issue better than GAP.
Figure 5. Comparison between $\mathsf{CARIBOU}$ (colored boxes) and GAP (blue lines) for ablation study on Chain-S.
Impact of $\epsilon$.
@@ 6.3. Ablation Study
We present the full results in Figure 9 in Appendix and one setting in Figure 5(b) ($K=10$ on Chain-S). Figure 7 in Appendix shows the accuracy of non-private $\mathsf{CARIBOU}$ and GAP(ϵ=infitalic-$\epsilon=\inf$) for comparison. When $K$ is much less than the number of nodes per chain, e.g., $K=1$ vs. 8-nodes chain, $\mathsf{CARIBOU}$ and GAP cannot realize satisfactory accuracy (both under 75%) even for non-private settings, as features from distant nodes cannot be effectively learnt. If $K$ is near to or larger than the number of nodes per chain (e.g., when $K=10$ for Chain-S, as shown in Figure 5(b)), though both GAP and $\mathsf{CARIBOU}$ see very high accuracy for non-private mode, the accuracy of GAP drops to 50% at $\epsilon=1$ and further decreases with increased $\epsilon$, suggesting the noise magnitude are not properly controlled. On the other hand, $\mathsf{CARIBOU}$ sees steady growth of accuracy along with increased $\epsilon$, which is a desired outcome for privacy protection. Notably, the accuracy of $\mathsf{CARIBOU}$ and GAP is lower than the Chain-S result in Table 3, because $K=6,7$ are chosen for best cases.
Finding 3.
@@
Private GNNs face fundamental trade-offs between privacy, utility, and model depth $K$. Chain-structured datasets can serve as an effective benchmark for long-range scenarios due to configurable topology. Model utility becomes more susceptible to the DP noise if GNNs are tightly coupled with the underlying graph structure. Stronger privacy requires larger noise, while deeper GNN architectures amplify noise effects although privacy converges as $K\rightarrow\infty$ theoretically.
Other hyper-parameters.
@@
Here, we evaluated the impact of other hyper-parameters in CGL, including $C_{\mathsf{L}},\alpha_{1},\beta$. For the ease of reading, we draw Figure 6 of classification accuracy using Chain-Sand Cora. As $C_{\mathsf{L}}$ constrains the features learned at each aggregation, in the relatively weak privacy guarantee ($\epsilon=16,32$), Figures 6(a) and 6(b) empirically confirms that the accuracy improves with $C_{\mathsf{L}}$ increases. In contrast, for strong privacy guarantee ($\epsilon=1,2$), larger $C_{\mathsf{L}}$ reduces the model accuracy due to the accumulated large noise. Figures 6(c) and 6(d) describe the ratio ($\alpha_{1}$) of learning from the graph, where $\alpha_{1}=1.0(\alpha_{2}=0.0)$ means the information from adjacent matrix is utilized at the maximum degree. Larger $\alpha_{1}$ leads to higher accuracy across varying $\epsilon$ in general, suggesting $\mathsf{CARIBOU}$ is able to achieve good balance between graph connectivity and privacy. The impact of $\beta$, which decides the power of residual connection between node features and CGL, is different on the two datasets. Since Chain-S is designed to tailor graph topology over node features, increasing $\beta$ to a large value (e.g., 15) might hurt accuracy. For Cora with rich node features, the model accuracy is generally increased along with $\beta$.
Figure 6. Classification accuracy under $C_{\mathsf{L}},\alpha_{1},\beta$ of CGL.
Finding 4.
All parameters $\sigma,C_{\mathsf{L}},\beta,\alpha_{1}$ in perturbed CGL contribute to the privacy-utility trade-off. Empirical analysis of $C_{\mathsf{L}},\beta,\alpha_{1}$ confirms the trade-offs by controlling the ratio of learnable node embeddings and initial node features, and topological attributes that affect the noise susceptibility.
7. Conclusion
@@
In this study, we provide a theoretical analysis establishing a convergent privacy budget for private GNNs. Our analysis addresses a longstanding limitation in perturbed message-passing architectures — namely, the linear accumulation of noise with depth — by showing that privacy loss can remain bounded as the number of layers increases. Consequently, deeper models can be deployed with a significantly improved privacy-utility trade-off. Our analysis is broadly applicable, requiring only two conditions that are commonly satisfied in practice: the use of hidden intermediate states — a standard design choice — and contractive message passing layers, which are often observed empirically.
@@
Appendix A Review on Message Passing GNNs
Every layer in the Graph Convolutional Network (GCN) (Kipf and Welling, 2017) can be expressed as a message passing layer, where the aggregation function simply computes weighted sums of the features of the neighbors. The GCN layer applied to graph ${\bm{G}}$ with a node feature matrix ${\bm{X}}^{(k)}$ can be expressed as:
(9) $\text{GCN}_{G}({\bm{X}}_{u}^{(k)})=\sigma\left(\hat{{\bm{A}}}{\bm{X}}^{(k)}{\bm{W}}^{(k)}\right),$
@@ Appendix A Review on Message Passing GNNs
Appendix B Preliminary on $f$-DP and Related Definitions
@@
The $f$-differential privacy framework (Dong et al., 2022; Kairouz et al., 2017) is based on hypothesis testing where $f$ denotes a trade-off function between type I and type II errors. Given two output distributions $P$ and $Q$ of a mechanism $\mathcal{M}$ on neighboring datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$, the problem is to distinguish between them: $H_{0}:\text{dataset is }\mathcal{D}$ vs. $H_{1}:\text{dataset is }\mathcal{D}^{\prime}$. For a rejection function $\phi:\mathcal{X}\to[0,1]$, the type I error is $\mathbb{E}_{P}\phi$ and the type II error is $1-\mathbb{E}_{Q}\phi$. The privacy guarantee is formalized through the trade-off function defined below.
@@
Definition 0 (Trade-off Function (Bok et al., 2024, Theorem 4.2)).
@@
A special case of $f$-DP is the Gaussian differential privacy (GDP) (Dong et al., 2022), which is defined as follows.
Definition 0 (Gaussian Differential Privacy (Bok et al., 2024, Definition 2.1)).
@@
A connivent tool for analyzing the composition of $f$-DP is the tensor product of trade-off functions, which is defined as follows.
Definition 0 ((Dong et al., 2022, Definition 3.1)).
@@
The tensor product of trade-off functions satisfies the following properties as proved in Dong et al. (2022, Proposition D.1).
Lemma 0 ((Dong et al., 2022, Proposition D.1)).
@@
Theorem 6 (GDP to RDP (Bok et al., 2024, Lemma A.4)).
@@
Theorem 7 (RDP to DP (Mironov, 2017, Proposition 3)).
@@
Appendix C Proofs for results in Section 5
C.1. Proof of results in Section 5.1

### v2

@@ VI-B Privacy Auditing
In Table XI (in Appendix), we report the AUC (Area Under the Curve) score about the attack effectiveness, when $\mathsf{CARIBOU}$ is being attacked. AUC is a major metric to evaluate the membership inference attack [sp/Wulinkteller22, he2021stealing]. Specifically, AUC measures true positive rate against the false positive rate on various classification thresholds, and a score of $0.5$ suggesting random guessing. We found $\mathsf{CARIBOU}$ is very effective against LinkTeller, by dropping the attack AUC from between 0.86 to 0.998 across all standard datasets ($\epsilon=\inf$) to less than 0.500 ($\epsilon$ ranges from 1 to 32). The similar effect was also observed by Wu et al. (e.g., less than 0.5 attack precision for 3-layer GCN for high density belief, shown in Table IX) [sp/Wulinkteller22] and Tang et al. (e.g., less than 0.5 attack AUC sometimes in Figure 10) [popets/TangNAA24]. For G-MIA, its AUC on the 5 datasets are already lower than LinkTeller by a notable margin when $\epsilon=\inf$ (between 0.567 to 0.702 for the 5 datasets), so the impact of $\mathsf{CARIBOU}$ is relatively small. But we observe on Cora, $\mathsf{CARIBOU}$ is able to drop AUC from 0.645 to 0.500.
Takeaway 2.
@@
VI-C Ablation Study
VI-C1 Impact of $K$
Both $\mathsf{CARIBOU}$ and GAP perform $K$-hop aggregations under $K$ aggregation layers. Here we evaluate the impact of $K$ on accuracy on the chain-structured datasets (Chain-S, Chain-M and Chain-L), as their classification results highly depend on long-range interactions. In Figure 9 and Figure 10 of Appendix, we compare $\mathsf{CARIBOU}$ and GAP on varying $\epsilon$ and varying $K$, respectively. The result of GAP is drawn with lines and the result of $\mathsf{CARIBOU}$ is illustrated with the colored boxes, because $\mathsf{CARIBOU}$ also depends on other hyper-parameters $C_{\mathsf{L}},\alpha_{1},\beta$ and we use the colored boxes to represent the interquartile range over 5 runs of their different value combinations.
In Figure 5(a), we show the result of one setting ($\epsilon=4$ on Chain-M), and $\mathsf{CARIBOU}$ achieves higher accuracy at every $K$. This observation is consistent in other settings as shown in Figure 9. In addition, the highest accuracy happens at $K=10$ (close to the number of nodes per chain) for $\mathsf{CARIBOU}$, and the classification accuracy fluctuates when $K$ varies for both $\mathsf{CARIBOU}$ and GAP. Across all datasets, GAP’s accuracy degrades monotonically with depth, consistent with its privacy noise variance growing linearly in the number of layers ($\sigma^{2}\propto K$). In contrast, $\mathsf{CARIBOU}$ benefits from additional depth and then plateaus, owing to its convergent privacy cost with respect to depth. This behavior shows that $\mathsf{CARIBOU}$ can leverage deeper architectures and realize the contractive privacy amplification guaranteed by our analysis.
Figure 5: Comparison between $\mathsf{CARIBOU}$ (colored boxes) and GAP (blue lines) for ablation study.
VI-C2 Impact of $\epsilon$
@@
We present the full results in Figure 10 in Appendix and one setting in Figure 5(b) ($K=10$ on Chain-S). Figure 7 in Appendix shows their non-private versions. When $K$ is much less than the number of nodes per chain, e.g., $K=1$ vs. 8-nodes chain, $\mathsf{CARIBOU}$ and GAP cannot realize satisfactory accuracy (both under 75%) even for non-private settings, as features from distant nodes cannot be effectively learnt. If $K$ is near to or larger than the number of nodes per chain (e.g., when $K=10$ for Chain-S, as shown in Figure 5(b)), though both GAP and $\mathsf{CARIBOU}$ see very high accuracy for non-private mode, the accuracy of GAP drops to 50% at $\epsilon=1$ and further decreases with increased $\epsilon$, suggesting the noise magnitude are not properly controlled. On the other hand, $\mathsf{CARIBOU}$ sees steady growth of accuracy along with increased $\epsilon$, which is a desired outcome for privacy protection.
Takeaway 3.
@@
Private GNNs face fundamental trade-offs between privacy, utility, and model depth $K$. Model utility becomes more susceptible to the DP noise if GNNs are tightly coupled with the underlying graph structure.
VI-C3 Noise scaling with depth
Here we analyze how the noise scale changes with the number of layers $K$ under standard linear composition (Corollary 2) and under convergent privacy analysis (Theorem 3). For a general study, we remove the effect of node degree $D_{\min},D_{\max}$ derived from a particular dataset and fix the target DP parameters $\epsilon,\delta$. In this case, the dependence of the calibrated noise on depth is governed by depth $K$. To make this comparison concrete, we instantiate a representative setting by normalizing the sensitivity: $\Delta^{2}(\mathrm{MP})=1,\alpha=6,C_{\mathsf{L}}=0.9.$ Setting $\Delta^{2}(\mathrm{MP})=1$ removes a common multiplicative factor and highlights the qualitative dependence on $K$. The choice $\alpha=6$ is common and simplifies the expressions, while $C_{\mathsf{L}}=0.9$ represents a standard contractive layer.
TABLE IV: Noise Scale $\sigma$ under Different $K$. We set $\Delta^{2}(\mathrm{MP})=1,\alpha=6,C_{\mathsf{L}}=0.9$ and $\epsilon=4,\delta=0.001$.
$K$ 11 22 44 88 $16$ $32$ $64$ $128$
Linear $1.07$ $1.52$ $2.15$ $3.04$ $4.30$ $6.07$ $8.56$ $12.11$
Convergent $1.07$ $1.52$ $2.14$ $3.00$ $4.07$ $4.66$ $4.66$ $4.67$
Table IV reports the proportional values of $\sigma$ for several representative depths $K$, assuming the same target privacy budget $\epsilon=4$. Under linear composition, the required $\sigma$ grows proportionally to $K$, becoming large for deep GNNs. In contrast, under $\mathsf{CARIBOU}$’s analysis, the required $\sigma$ grows from $1$ to a bounded constant (here approximately $4.7$) and then saturates. This study highlights $\mathsf{CARIBOU}$ can support deep architectures without unbounded noise growth. In addition, a failure case of divergent noise allocation has been shown in Appendix A-H.
VI-C4 hyper-parameters related to contractiveness
We studied the impact of hyper-parameters $C_{\mathsf{L}},\alpha_{1},\beta$ in CGL. In Appendix, we draw Figure 11 of classification accuracy using Chain-Sand Cora. As $C_{\mathsf{L}}$ constrains the features learned at each aggregation, in the relatively weak privacy guarantee ($\epsilon=16,32$), Figures 11(a) and 11(b) empirically confirms that the accuracy improves with $C_{\mathsf{L}}$ increases. In contrast, for strong privacy guarantee ($\epsilon=1,2$), larger $C_{\mathsf{L}}$ reduces the model accuracy due to the accumulated large noise. Small $C_{\mathsf{L}}$ enforces strong contraction, accelerating privacy convergence and reducing effective sensitivity, but overly small values may reduce expressive power. Larger $C_{\mathsf{L}}$ increases representational capacity but slows contraction and slightly increases noise amplification.
@@ VI-C3 Noise scaling with depth
Figures 11(c) and 11(d) describe the ratio ($\alpha_{1}$) of learning from the graph, where $\alpha_{1}=1.0(\alpha_{2}=0.0)$ means the information from adjacent matrix is utilized at the maximum degree. Larger $\alpha_{1}$ leads to higher accuracy across varying $\epsilon$ in general, suggesting $\mathsf{CARIBOU}$ is able to achieve good balance between graph connectivity and privacy. The impact of $\beta$, which decides the power of residual connection between node features and CGL, is different on the two datasets. Since Chain-S is designed to tailor graph topology over node features, increasing $\beta$ to a large value (e.g., 15) might hurt accuracy. For Cora with rich node features, the model accuracy is generally increased along with $\beta$.
Takeaway 4.
All parameters $\sigma,C_{\mathsf{L}},\beta,\alpha_{1}$ in perturbed CGL contribute to the privacy-utility trade-off.
VI-C5 Impact of $D_{\max},D_{\min}$
Figure 6 shows an example ($\epsilon=2$) of the classification accuracy of NDP under different maximum node degrees, and more ablation study results are deferred to Figure 8. As shown in Figure 6(a) on the Photo dataset, $\mathsf{CARIBOU}$ consistently realizes the highest accuracy under different maximum node degrees ranging from $5$ to $100$. Improving maximum node degree for DPDGC can slightly increase the classification accuracy when maximum node degree is $20$, while the number of maximum node degree does not help for PertGraph and GAP. For the Chain-S dataset in Figure 6(b), classification accuracy of $\mathsf{CARIBOU}$ is improved when the maximum node degree is increased and relatively small. In addition, $\mathsf{CARIBOU}$ outperforms baseline works significantly, i.e., approximated $20\%$-$25\%$ higher than the second best GAP.
The sensitivity formulas in Theorems 4,5 explicitly depend on structural properties of the graph, particularly minimum degree $D_{\min}$ and maximum degree $D_{\max}$. Our empirical results (Table III, Figure 8) reflect these theoretical dependencies: (i) High-degree graphs such as Photo exhibit lower noise and higher accuracy; (ii) Low-degree or chain-like graphs incur higher sensitivity and lower accuracy, but $\mathsf{CARIBOU}$ mitigates the impact.
Figure 6: NDP Accuracy with Varying Max Node Degree.
VII Future Works and Discussions
More GNN models. $\mathsf{CARIBOU}$ is instantiated and evaluated primarily with commonly used message-passing architectures, where contractiveness naturally emerges or can be enforced by design. Extending our convergent privacy framework to a broader class of GNNs, including attention-based models (e.g., GAT), spectral convolution methods, and emerging graph transformers would be a natural next step, but new research problems will emerge. For instance, these models differ in their Lipschitz properties, aggregation operators, and feature mixing patterns, which may influence the achievable privacy amplification and the expressiveness–contractiveness trade-off. Developing a unified analysis for these graph families, or designing contractive variants of non-message-passing architectures, is an open and promising direction.
White-box attacks and defenses. Our privacy auditing focuses on black-box membership threats, which align with the theoretical guarantees of perturbed message passing. However, stronger adversaries with white-box access to gradients, intermediate embeddings, or partial training states can mount reconstruction, inversion, or property inference attacks that fall outside our current threat model. Prior work has shown that gradient-based attacks can recover fine-grained structural information, especially in over-parameterized models. Investigating the extent to which contractiveness mitigates these stronger threats, and designing DP mechanisms that remain robust under partial or full white-box exposure, warrant future research. Such analyses may require combining $\mathsf{CARIBOU}$ with complementary techniques such as gradient perturbation, secure aggregation, or private feature compression.
VIII Related Works
Multi-layer GNNs. Recent literature shows that multi-layer GNNs hold significant potential for modeling long-range dependencies and complex relational structures crucial for many real-world applications. Node labels and attributes may depend on distant nodes, necessitating the aggregation of information over larger receptive fields [tnn/WuPCLZY21] through multi-layer GNNs. Notably, Li et al. [icml/Li0GK21] demonstrated, through the 1000-layer GNN, that increasing the network depth attains substantial gains in accuracy, e.g., from $72\%$ with shallow GNNs to $88\%$ with hundred- and thousand-layer GNNs, by capturing distant features. However, enforcing DP in multi-layer GNNs is particularly challenging, as these GNN models aggregate node embeddings over deeper layers and broader neighborhoods. Current research still lacks an effective solution to injecting DP noise to multi-layer GNNs with privacy-utility balance.
Differentially private GNNs. Graphs consist of edges and nodes. Corresponding to instance-level DP [csf/mironov2017renyi, ccs/AbadiCGMMT016, nips/0001S22], the “instance” of graphs can be an edge or a node, naturally called edge-level DP (EDP) and node-level DP (NDP). GNNs have emerged as a key approach for applications over graph-structured data, such as intrusion detection [eurosp/ApruzzeseLS23, sp/SommerP10], social recommendation [icde/XiaSH0XP23], and drug discovery [kdd/ZhongM23]. Sharing trained GNN model can lead to privacy risks [uss/000100S022, ccs/0001MMBS22, corr/abs-2102-05429], typically membership inference attack (MIA) [sp/CarliniCN0TT22, ccs/0001MMBS22]. MIA stems from “overfitting”, where models can memorize training memberships [zhang2021understanding], either an edge or a node. Consequently, GNNs can leak sensitive information about their edge- or node-level neighbors.
To address these risks, existing research works [arxiv/sajadmanesh2020differential, arxiv/Wuprivacy23, nips/chien2023differentially, arxiv/DaigavaneNode21, corr/abs-2109-08907, uss/sajadmanesh2023gap] have integrating DP with GNNs to achieve EDP and NDP. One research direction is to utilize graph perturbation (e.g., LPGNet [ccs/KolluriBHS22] and LapGraph [sp/0011L0022]) through a randomized response mechanism and adding discrete DP noise to the adjacency matrix. Then, the perturbed graph is passed to GNNs for subsequent training tasks, where the graph perturbation is required only once and also irrelevant to the GNN architectures. However, the GNN model utility is low when being trained over a perturbed graph when the privacy budget is tight, for example, $40\%$ accuracy of $\epsilon=1,2,3,4$ reported in LPGNet [ccs/KolluriBHS22].
To improve utility, perturbed message-passing mechanism (PMP) [uss/sajadmanesh2023gap] has been proposed by adding the calibrated Gaussian noise to the message-passing layer, and DPDGC perturbs the decoupled graph convolution [nips/chien2023differentially]. As PMP realizes a better trade-off of privacy and utility, our work extends the research line of PMP. Table I presents a comprehensive comparison. Albeit their efforts on EDP and NDP, leveraging the contractive hidden node embeddings in private GNNs for amplifying privacy remains an underexplored avenue; thus, $\mathsf{CARIBOU}$ fills this gap. More related works are detailed in Appendix A-I.
IX Conclusion
@@
In this study, we provide a theoretical analysis establishing a convergent privacy budget for private deeper GNNs. Our analysis addresses a longstanding limitation in perturbed message-passing architectures, namely, the linear accumulation of noise with depth, by showing that privacy loss can remain bounded as the number of layers increases. Consequently, deeper models can be deployed with a significantly improved privacy-utility trade-off. Our analysis is broadly applicable, requiring only two conditions that are commonly satisfied in practice: the use of hidden intermediate states (also a standard design choice) and contractive message passing layers, which are often observed empirically.
@@
Acknowledgment
We want to thank the reviewers for their insightful and constructive comments. The authors of University of California, Irvine are partially supported by NSF CNS-2220434 and TIP-2453148. Any opinions, findings, and conclusions expressed in this material are those of the author(s) and do not necessarily reflect the views of the sponsors.
-A Review on Message Passing GNNs
Every layer in the Graph Convolutional Network (GCN) [iclr/kipf2017semi] can be expressed as a message passing layer, where the aggregation function simply computes weighted sums of the features of the neighbors. The GCN layer applied to graph $\bm{G}$} with a node feature matrix ${\bm{X}}^{(k)}$ can be expressed as:
$\text{GCN}_{G}({\bm{X}}_{u}^{(k)})=\sigma\left(\hat{{\bm{A}}}{\bm{X}}^{(k)}{\bm{W}}^{(k)}\right),$ (9)
@@ Acknowledgment
Appendix A Preliminary on $f$-DP and Related Definitions
@@
The $f$-differential privacy framework [journal/dong2022gaussian, tit/KairouzOV17] is based on hypothesis testing where $f$ denotes a trade-off function between type I and type II errors. Given two output distributions $P$ and $Q$ of a mechanism $\mathcal{M}$ on neighboring datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$, the problem is to distinguish between them: $H_{0}:\text{dataset is }\mathcal{D}$ vs. $H_{1}:\text{dataset is }\mathcal{D}^{\prime}$. For a rejection function $\phi:\mathcal{X}\to[0,1]$, the type I error is $\mathbb{E}_{P}\phi$ and the type II error is $1-\mathbb{E}_{Q}\phi$. The privacy guarantee is formalized through the trade-off function defined below.
@@
Definition 9 (Trade-off Function [icml/BokShifted24, Theorem 4.2]).
@@
A special case of $f$-DP is the Gaussian differential privacy (GDP) [journal/dong2022gaussian], which is defined as follows.
Definition 10 (Gaussian Differential Privacy [icml/BokShifted24, Definition 2.1]).
@@
A common tool for analyzing the composition of $f$-DP is the tensor product of trade-off functions, which is defined as follows.
Definition 11 ([journal/dong2022gaussian, Definition 3.1]).
@@
The tensor product of trade-off functions satisfies the following properties as proved in [journal/dong2022gaussian, Proposition D.1].
Lemma 2 ([journal/dong2022gaussian, Proposition D.1]).
@@
Theorem 7 (GDP to RDP [icml/BokShifted24, Lemma A.4]).
@@
Theorem 8 (RDP to DP [csf/mironov2017renyi, Proposition 3]).
@@
A-A Preliminaries: Perturbed Message Passing
Recall that each layer of a perturbed message-passing GNN updates the node features following Equation 2, where ${\bm{Z}}\sim\mathcal{N}\bigl(0,\sigma^{2}\mathrm{I}_{d}\bigr)$ is Gaussian noise added to obscure the true output of $\mathrm{MP}_{{\bm{G}}},$ and $\Pi_{\mathcal{K}}$ projects onto some convex feasible set $\mathcal{K}$}. As shown in Proposition 2, for a single layer, the privacy cost depends only on $\Delta(\mathrm{MP})$ and $\sigma.$ However, by the composition theorem of RDP [csf/mironov2017renyi], stacking $K$ such layers yields a cost that scales linearly in $K$ (Corollary 2).
Proposition 2.
[…]
Corollary 2.
By applying the composition theorem of Rényi differential privacy, a $K\text{-layer}$ perturbed message-passing GNN satisfies $\Bigl(\alpha,\frac{K\,\alpha\,\Delta^{2}(\mathrm{MP})}{2\sigma^{2}}\Bigr)\text{-RDP},$ which implies $\Bigl(\frac{K\,\alpha\,\Delta^{2}(\mathrm{MP})}{2\sigma^{2}}+\frac{\log(1/\delta)}{\alpha-1},\delta\Bigr)\text{-DP}.$
This linear growth of the privacy budget is common in multi-hop settings, such as [uss/sajadmanesh2023gap], but is troublesome for deep GNNs, especially when many message-passing steps are needed for long-range dependencies. Excessively large privacy loss often undermines the model’s utility.